The historical returns on a balanced portfolio have had an average return of 13% and a standard deviation of 16%. Assume that returns on this portfolio follow a normal distribution. [You may find it useful to reference the z table.]

a. What percentage of returns were greater than 45%? (Round your answer to 2 decimal places.)



b. What percentage of returns were below −35%? (Round your answer to 2 decimal places.)
Answer #1

Solution :

Given that ,

mean = = 13% = 0.13

standard deviation = = 16% = 0.16

a)

P(x > 45%) = 1 - P(x < 0.45)

= 1 - P((x - ) / < (0.45-0.13) / 0.16)

= 1 - P(z <2 )

= 1 - 0.9772   

= 0.0228

= 2.28%

Percetage = 2.28%

b)

P(x <-35% ) = P((x - ) / < (-0.35-0.13) /0.16 )

= P(z <-3 )

= 0.0013

= 0.13%

Percentage = 0.13%